a rectangle is 3 times aas long as it is wide. if the lenght is increased by 6 and the width by 8 the area is increased by 109 find the original dimensions
Answer by richard1234(7193)

Suppose the dimensions of the original rectangle are x and 3x. Then

(3x + 6)(x + 8) = 3x^2 + 109

3x^2 + 30x + 48 = 3x^2 + 109

30x + 48 = 109

x = 61/30, 3x = 61/10